# Point-PEFT: Parameter-Efficient Fine-Tuning for 3D Pre-trained Models

### 06 Dec 2023 v4

@@ Abstract
The popularity of pre-trained large models has revolutionized downstream tasks across diverse fields, such as language, vision, and multi-modality. To minimize the adaption cost for downstream tasks, many Parameter-Efficient Fine-Tuning (PEFT) techniques are proposed for language and 2D image pre-trained models. However, the specialized PEFT method for 3D pre-trained models is still under-explored. To this end, we introduce Point-PEFT, a novel framework for adapting point cloud pre-trained models with minimal learnable parameters. Specifically, for a pre-trained 3D model, we freeze most of its parameters, and only tune the newly added PEFT modules on downstream tasks, which consist of a Point-prior Prompt and a Geometry-aware Adapter. The Point-prior Prompt adopts a set of learnable prompt tokens, for which we propose to construct a memory bank with domain-specific knowledge, and utilize a parameter-free attention to enhance the prompt tokens. The Geometry-aware Adapter aims to aggregate point cloud features within spatial neighborhoods to capture fine-grained geometric information through local interactions. Extensive experiments indicate that our Point-PEFT can achieve better performance than the full fine-tuning on various downstream tasks, while using only 5% of the trainable parameters, demonstrating the efficiency and effectiveness of our approach. Code will be released at https://github.com/Even-JK/PEFT-3D.

### 16 Dec 2023 v5

@@ Abstract
The popularity of pre-trained large models has revolutionized downstream tasks across diverse fields, such as language, vision, and multi-modality. To minimize the adaption cost for downstream tasks, many Parameter-Efficient Fine-Tuning (PEFT) techniques are proposed for language and 2D image pre-trained models. However, the specialized PEFT method for 3D pre-trained models is still under-explored. To this end, we introduce Point-PEFT, a novel framework for adapting point cloud pre-trained models with minimal learnable parameters. Specifically, for a pre-trained 3D model, we freeze most of its parameters, and only tune the newly added PEFT modules on downstream tasks, which consist of a Point-prior Prompt and a Geometry-aware Adapter. The Point-prior Prompt adopts a set of learnable prompt tokens, for which we propose to construct a memory bank with domain-specific knowledge, and utilize a parameter-free attention to enhance the prompt tokens. The Geometry-aware Adapter aims to aggregate point cloud features within spatial neighborhoods to capture fine-grained geometric information through local interactions. Extensive experiments indicate that our Point-PEFT can achieve better performance than the full fine-tuning on various downstream tasks, while using only 5% of the trainable parameters, demonstrating the efficiency and effectiveness of our approach. Code is released at https://github.com/Ivan-Tang-3D/PEFT-3D.
@@ 7 Conclusion
Appendix A Implementation Details
The inner dimension of the bottleneck-shaped MLP in the Point-prior Prompt is set as 8 for all pre-trained models on two datasets. The inner dimension of the first bottleneck-shaped MLP in the Geometry-aware Adapter is decided as 16 for all pre-tarined mdoels. We also set different N and k for the k-NN algorithm at different pre-trained models. The N and k are set as (32, 16), (32, 8), and (64, 16) for Point-BERT [28], Point-M2AE [32], and Point-MAE [17] respectively. On the ScanObjectNN [24] and ModelNet40 [25], we set the inner dimension of the last bottleneck-shaped MLP in the Adapter as (8, 16, 16) and (16,16, 16) for Point-BERT, Point-M2AE, and Point-MAE respectively.
Methods #Param(M) Acc.(%)
Training from Scratch
Point-NN 0.0 64.9
PointNet 3.5 68.0
PointNet++ 1.5 77.9
PointMLP 14.9 85.2
${}^{\dagger}$ 0.8 87.1
Self-supervised Pre-training
Point-BERT 22.1 83.1
w/ Point-PEFT 0.6 85.0
$\downarrow$97.4+1.9
Point-M2AE 15.3 86.4
w/ Point-PEFT 0.7 86.4
$\downarrow$95.4+0.0
${}^{\dagger}$ 15.3 88.1
w/ ${}^{\dagger}$ 0.7 88.2
$\downarrow$95.4+0.1
Point-MAE 22.1 85.2
w/ Point-PEFT 0.7 85.5
$\downarrow$96.8+0.3
${}^{\dagger}$ 22.1 88.1
w/ ${}^{\dagger}$ 0.7 89.1
$\downarrow$96.8+1.0
Table 7: Performance of Shape Classification on ScanObjectNN Dataset. We report the number of learnable parameters (#Param) and the accuracy (%) on the ”PB-T50-RS” split of ScanObjectNN [24]. $\dagger$ indicates utilizing a stronger data augmentation [37] during fine-tuning.
Appendix B Additional Experiments
Shape Classification.
Due to the page limitation, we only provide part of the results of Point-PEFT for 3D shape classification in the main paper. In Table 7, we report more quantitative results of Point-PEFT with the stronger data augmentation (random scaling, translation and rotation) and default augmentation method (random scaling and translation), which are denoted as with and without $\dagger$ respectively. As reported, under both of the two data augmentation settings, our Point-PEFT outperforms the full fine-tuning method based on almost all pre-trained models. This indicates the robustness of Point-PEFT on different pre-trained 3D point cloud models and data augmentation methods.
[…]
Training from Scratch
PointNet - 80.39 83.7
PointNet++ - 81.85 85.1
DGCNN - 82.33 85.2
Self-supervised Pre-training
Transformer 27.09 83.42 85.1
MaskPoint - 84.60 86.0
${}^{\ddagger}$ 22.26 84.19 86.1
w/ Point-PEFT 0.45 83.81 85.3
${}^{\ddagger}$ 13.03 84.86 86.5
w/ Point-PEFT 0.39 84.35 86.1
[…]
Part Segmentation.
[…]
         Positions of GA-Adapter. Insertion Depth Acc.(%)
    After    Before   Both   Parallel
- - - 6 88.0%
- - - 12 89.1%
- - - 6 88.3%
- - - 12 88.0%
- - - 6 86.0%
- - - 12 86.1%
- - - 6 87.9%
- - - 12 88.4%
Table 9: Ablation Study on Adapter Positions and Depth. The ‘After’, ‘Before’, ‘Both’, ’Parallel’ denote block structure of putting the ”Geometry-aware Adapter” after, before, wrapping around, and in parallel to the ”FFN” layer respectively. The insertion depth denotes the certain number of earlier blocks.
Appendix C Additional Ablation Study
The Adapter Positions and Depth
We further conduct ablation study on the positions and depth of the proposed Geometry-aware Adapter in Table 9, where we utilize the final version of Point-PEFT as the baseline and compare the accuracy(%) on the ”PB-T50-RS” split of the ScanObjectNN with ${}^{\dagger}$ as the pre-trained model. As shown in Table 9, the last position within the block and the depth of 12 perform the best. That is because, compared to other positions, after the global interactions among point cloud features through pre-trained self-attention layer and Feed-forward Networks(FFN), the local interactions within the adapter could further boost the extraction of the fine-grained geometric information, which facilitates the fine-tuning process. And the deeper insertion could utilize the high-level knowledge to grasp the better understanding of 3D structures. Note that, for all pre-trained point cloud models, we insert the Geometry-aware Adapter in the last position across all blocks.